OpenStudy (anonymous):

The following is a parallelogram. Find x and y using the given information. AB=5x-2y DC=4 AD=y BC=5-x

OpenStudy (anonymous):

diagram ?

OpenStudy (anonymous):

use the lengths of the opposite sides are equal for parallelogram

OpenStudy (anonymous):

or i guess x is 2 and y is 3

OpenStudy (anonymous):

for guessing i have assumed 5x-2y=4 or AB=DC and 5-x=y or AD=BC

OpenStudy (anonymous):

AB = DC 5x - 2y = 4 AD = BC y = 5 - x substitute second equation into the first 5x - 2(5 - x) = 4 simplify 5x - 10 + 2x = 4 combine like terms 7x - 10 = 4 add 10 to both sides 7x -10 + 10 = 4 + 10 simplify 7x + 0 = 14 7x = 14 divide by 7 7x/7 = 14/7 simplify 1x/1 = 2 x = 2 back substitute into either of the original two equations to get y. 5(2) - 2y = 4 10 - 2y = 4 10 - 4 = 2y 6 = 2y 3 = y conclude that x = 2 and y = 3.